. Find the area included between
x^2+y^2= 4x
and
y^2=x
above x-axis
1
Expert's answer
2021-09-29T00:51:28-0400


Find point of intersection:

"x^2=3x\implies x=3,y=\sqrt{3}"


"S=S_1+S_2"

where "S_1" is area under the graph of "x=y^2" ,

"S_2" is area under the graph of "x^2+y^2=4x" at "x\isin [3,4]"


"S_1=\int^3_0 \sqrt{x}dx=\frac{2x\sqrt{x}}{3}|^3_0=2\sqrt{3}"


S2 is half of area of sector with central angle "\pi/3" minus area of triangle with vertices: (2,0),(3,0),(3,"\sqrt{3}" ).

"S_2=R^2\alpha/4-\sqrt{3}/2"

where "R=2" ,

"\alpha=tan^{-1}(\sqrt{3})=\pi/3"

"S_2=4\pi/12-\sqrt{3}/2=\pi/3-\sqrt{3}/2"


"S=2\sqrt{3}+\pi/3-\sqrt{3}/2=\pi/3+1.5\sqrt{3}=3.645"